OpenStudy (haleyelizabeth2017):

Rewrite ³√27x - 81 - 5 to make it easy to graph using a translation. Describe the graph. I am not sure where to start.......

OpenStudy (haleyelizabeth2017):

\[\sqrt[3]{27x-81}-5\]

OpenStudy (haleyelizabeth2017):

@confluxepic are you available right now to help me out? lol I can't remember for the life of me haha yeah....

OpenStudy (confluxepic):

Hold on. Oregon teaches different things than Texas. I am learning this right now. I will be back in 5 minutes with full understanding of this specific question.

OpenStudy (haleyelizabeth2017):

okay thank you so much!

OpenStudy (confluxepic):

First you do -81-5 so it will become -86.

OpenStudy (haleyelizabeth2017):

but wouldn't it have to be the cube root of -81 first? because the cube root is for both the 27x as well as the -81.......This stuff if confusing as heck :/

OpenStudy (confluxepic):

Cubed root of 27 would be 3 but I am not sue what to do with x. And yes. Cubed rot of -81 is better.

OpenStudy (confluxepic):

But you can't root a negative number.

OpenStudy (haleyelizabeth2017):

true....

OpenStudy (confluxepic):

Then i will have to come in. \[i=\sqrt{-1}\]

OpenStudy (haleyelizabeth2017):

okie dokie........I am not liking math right now..... :/ just plain confusing

OpenStudy (confluxepic):

@TheSmartOne know more about i.

OpenStudy (haleyelizabeth2017):

okay

OpenStudy (confluxepic):

https://www.desmos.com/calculator/nwlqj0uz7u

OpenStudy (haleyelizabeth2017):

Sadly, I have to show work to rewrite it :(

OpenStudy (haleyelizabeth2017):

@satellite73 might be able to help

OpenStudy (confluxepic):

http://www.mathsisfun.com/numbers/cube-root.html

OpenStudy (confluxepic):

\[\sqrt[3]{-81}=-4.32674871092\]

OpenStudy (confluxepic):

@haleyelizabeth2017 We skipped out on one fact. It is cubed root not square root so it is possible for a negative number to be cubed rooted.

OpenStudy (haleyelizabeth2017):

okay......thats what I thought because it worked in my calc lol

OpenStudy (haleyelizabeth2017):

we need to factor....just looked it up lol

OpenStudy (haleyelizabeth2017):

so it will be \[3\sqrt[3]{(x-3)}-5\] and the graph is: -3 = Horizontal shift (To the right). -5 = Vertical shift (Down five units).

OpenStudy (confluxepic):

Oh. You factored it. Never thought about that.

OpenStudy (haleyelizabeth2017):

lol thank you for trying to help :)

OpenStudy (confluxepic):

You're welcome. I am sure we both learned from this.

OpenStudy (haleyelizabeth2017):

lol yeppers